# Multi Task Inverse Reinforcement Learning for Common Sense Reward

### v1

@@ 4.1 Curriculum Learning
To overcome this issue and avoid any serious degradation to the task-specific performance, we devised a simple curriculum learning approach (Bengio et al., 2009). Instead of updating the policy with the combined reward $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime})$, we use $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+\alpha(\bar{r}_{t}(s,a,s^{\prime}))f_{\theta}(s,a,s^{\prime})$ with the adaptive weighting $\alpha(\bar{r}_{t}(s,a,s^{\prime}))=\frac{\bar{r}_{t}(s,a,s^{\prime})}{R_{MAX}}$ assuming $\bar{r}_{t}(s,a,s^{\prime})\in[0,R_{MAX}]$.
This allows us to gradually increase the weighting of our learned reward as the policy improves on the main task. Additionally, the common-sense reward does not have a significant effect at the beginning of training, and thus only impacts the policy after it had several update rounds. We note that this does not affect the discriminator update, which still uses Eq. 6. Also, when training a new agent with our learned reward we train from scratch using the standard aggregation $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime})$.
reach-wall button-press- topdown-wall drawer-close push-back coffee-button
reach-wall $\mathbf{0.93}$ $0.40$ $0.43$ $0.40$ $0.25$
button-press-topdown-wall $0.36$ $\mathbf{0.82}$ $0.50$ $0.31$ $0.21$
drawer-close $0.34$ $0.48$ $\mathbf{0.89}$ $0.35$ $0.24$
push-back $0.25$ $0.25$ $0.20$ $\mathbf{0.88}$ $0.22$
coffee-button $0.28$ $0.32$ $0.35$ $0.38$ $\mathbf{0.91}$
Table 1: Velocity cs-reward. Each row represents a different task for MT-CSIRL training, each column represents a task for RL training and evaluation. The Numbers represent the ratio between the agents’ velocity and its target value.
@@ 5 Experiments
For our experiments, we use the Meta-world benchmark (Yu et al., 2019). It provides a diverse set of robotic manipulation tasks, which share the same robot, action space, and observation space. Thus, this benchmark is suitable for evaluating the effectiveness and transferability of our inverse reinforcement learning method across various scenarios. We trained all policies using the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018). In order to emphasize our primary goal of learning transferable rewards, we limit our experimental setup to Meta-world tasks on which SAC performed well, according to the Meta-world’s paper. We used the exact training process and SAC hyperparameters as in (Yu et al., 2019). We repeat the experiments five times using different random initializations and report the mean and standard deviation of the performance.
Using the Meta-world terminology, there are several distinct tasks, e.g., reach-wall and drawer-close. For each task, there are several variations with distinct targets. For example, in the drawer-close task, a different target would be a different location of the drawer. We also note that Meta-world has two versions for each task. All experiments in this paper use the second version, i.e., reach-wall is reach-wall-V2, and we omit the V2 for brevity.
@@
The goal of our experiments is to demonstrate that IRL fails to learn a useful cs-reward and that training the cs-reward on multiple tasks is the key to learning such a reward. To show that we perform several experiments, each time learning the cs-reward (which we will describe shortly) from a more diverse set of tasks. See Fig. 2 for a visual summery of our experiments.
5.1 Common-Sense Rewards
@@
As Meta-world only contains task-specific rewards, we designed two simple common-sense rewards for our experiments. Our first cs-reward directs the agent to move one of the key points along the robotic arm (whose 3D location is part of the observation vector) with a target velocity $v_{target}$. The reward function we use is
$r_{CS}(s,a,s^{\prime})=-C_{v}\cdot|\|\ell(s^{\prime})-\ell(s))\|_{2}-v_{target}|$ (8)
@@
where $\ell(s)$ is the 3D location of the selected key point given the observation vector. The second reward directs the $L_{1}$ norm of the action towards a target value $n_{target}$. The reward is defined as follows,
$r_{CS}(s,a,s^{\prime})=-C_{n}\cdot|\|a\|_{1}-n_{target}|$ (9)
We name these the velocity and action norm cs-rewards, respectively. See Appendix C.1 for further details.
@@
One important property of these rewards is that they do not depend on the task or environment, and as such should be easily transferable. We specifically designed these simple common-sense behaviors to show how IRL struggles to learn a transferable reward.
Figure 5: Scatter plot of learned cs-reward and ground-truth cs-reward in the single-task transfer experiment, with velocity as common sense.
reach-wall button-press-topdown-wall drawer-close push-back coffee-button
reach-wall $\mathbf{0.91}$ $0.43$ $0.60$ $0.54$ $0.58$
button-press-topdown-wall $0.60$ $\mathbf{0.91}$ $0.58$ $0.60$ $0.59$
drawer-close $0.59$ $0.53$ $\mathbf{0.90}$ $0.63$ $0.57$
push-back $0.56$ $0.56$ $0.51$ $\mathbf{0.90}$ $0.55$
coffee-button $0.58$ $0.57$ $0.54$ $0.58$ $\mathbf{0.92}$
Table 2: Action Norm cs-reward. Each row represents a different task for MT-CSIRL training, each column represents a task for RL training and evaluation. The Numbers represent the ratio between the agents’ action norm and its target value.
5.2 CS-Reward Trained on a Single Task
Here, we empirically demonstrate that, even for simple common-sense rewards for which the IRL process successfully trains an agent, the final reward cannot be used to train a new agent from scratch. This limitation persists even when applying the reward on the exact same task and target.
@@ 5.2 CS-Reward Trained on a Single Task
To that end, we train an expert with the task-specific (i.e., reach-wall) and ground truth common-sense reward for each of our common-sense rewards. In both cases, we successfully train experts who maximize both the task rewards and the common-sense rewards. Then, for each expert, we sample a set of 60K trajectories for the IRL process. The cs-reward is learned using Alg. 1 with a single task (i.e., $T=1$). Finally, we train a new agent on the exact same task and target. In all of the following experiments in this section, all trained agents achieved 100% success rate on the main task. Therefore, we only report the results on the common-sense rewards.
@@
We present our results in Figure 3. During the IRL process, the trained agent effectively acquired the desired common-sense behavior, marked in the horizontal red line. However, when we train a new RL agent with our learned reward (and the task-specific reward) does not achieve the same desired behavior, even when training on the exact same target. Similar results with different tasks are shown in Appendix A. We note that this phenomenon has been also previously observed in other bi-level optimization scenarios (Navon et al., 2020; Vicol et al., 2022).
@@
In Figure 5, we also present a scatter plot of the learned cs-reward versus the ground-truth cs-reward. This shows a very small correlation between the learned reward and the ground-truth reward (correlation coefficient of 0.18), which further demonstrates that the IRL fails to capture the desired reward.
Figure 6: Scatter plot of ground-truth VS learned cs-reward for MT-CSIRL on velocity experiment. The rewards are calculated on an unseen task.
5.3 Learning CS-Reward with Multiple Targets
In the next experiment we show that when we train our IRL on expert demonstration from different targets, we manage to learn a reward that can be used to train new agents on unseen targets. However, it does not transfer to novel tasks. Again, all agents trained in this section achieve 100% success rate so we only show results for the cs-reward.
@@ 5.3 Learning CS-Reward with Multiple Targets
To do that, we train our MT-CSIRL method with experts trained on different targets for the same task, for each of our cs-rewards. We then test how these learned rewards can be used to train on new targets in this environment and how they transfer to other tasks. We execute this experiment on five different Meta-world tasks: reach-wall, button-press-topdown-wall, drawer-close, push-back and coffee-button. For each task, we did the following: (i) Train an expert for each of the 5 different targets. (ii) Sampled 10K trajectories from each expert. (iii) Train cs-reward using MT-CSIRL. (iv) Train with the learned cs-reward and evaluate using both new targets and new tasks.
The results for both velocity and action norm are presented in Table 1 and 2, respectively. The numbers represent the ratio between the agents’ velocity or action norm and its target value, i.e., the closer to one, the better. On the diagonal, we see the results on different targets for the original task, and on the off-diagonal, we see the results when testing on new tasks.
@@
When looking at the diagonal elements of Tables 1 and 2, we see that our reward transfers well to new targets for the seen task. While our agents do not reach the target values, we still get relatively close to them. However, the drop in performance on the off-diagonal elements indicates that the learned reward does not transfer well to novel tasks. This is especially interesting when we consider our simple cs-rewards, which do not depend on the specifics of the environment (and in the action norm case, is entirely independent of the state).
Table 3: MT-CSIRL experiment results on unseen task with action norm cs-reward.
Agent Avg Action Norm Success-rate
Expert $2.95\pm 0.01$ $100\pm 0.00$
SAC $1.57\pm 0.48$ $100\pm 0.00$
MT-AIRL $1.30\pm 0.47$ $63.5\pm 1.20$
MT-CSIRL $2.75\pm 0.18$ $99.9\pm 0.01$
Table 4: MT-CSIRL experiment results on unseen task with velocity cs-reward.
Agent Avg velocity Success-rate
Expert $0.81\pm 0.02$ $100\pm 0.00$
SAC $0.32\pm 0.16$ $100\pm 0.00$
MT-AIRL $0.11\pm 0.01$ $67.5\pm 5.90$
MT-CSIRL $0.69\pm 0.07$ $98.9\pm 0.02$
5.4 Learning CS-Reward with Multiple Tasks
@@
In our final experiment, we show that when we train the IRL process on experts’ trajectories from multiple tasks, we learn a useful cs-reward that can be transferred to new unseen tasks. In order to do that, we perform our MT-CSIRL method again, but this time we train each expert on a different meta-world task, as shown in Figure 2.
We evaluate the following methods: (i) Expert: RL trained with the task and ground-truth cs-reward (ii) SAC: An RL trained with only the task reward (iii) MT-AIRL: For this baseline, we employ a multi-task AIRL for learning the entire combined reward per task. This baseline also receives expert trajectories from the test task. (iv) MT-CSIRL: Our proposed method. An RL agent trained with the task and learned cs-reward on an unseen task.
@@
The results for velocity as cs-reward, and action norm as cs-reward are presented in table 3 and 4, respectively. As can be easily observed, the MT-CSIRL results show that our learned cs-reward manages to transfer the desired behavior even to unseen tasks, albeit not as strongly as the ground-truth reward. We included the MT-AIRL, which unsurprisingly does not work well, to show the importance of our split between task-specific and task-independent rewards. Without this distinction combining different tasks is non-trivial and can easily harm performance as we see here. We also note that our MT-AIRL baseline shares similarities with (Gleave & Habryka, 2018), however, they train with a meta-learning approach similar to MAML.
@@
To further illustrate our results we show in Fig. 4 the ground truth cs-reward (both for velocity and action norm) on the unseen test task. The figures show the positive impact of our cs-reward, making the agent’s behavior significantly more aligned with the expert. Finally, we show in Fig. 6 a scatter plot of the ground-truth cs-reward versus our learned reward on a new unseen task. As one can see, these rewards are highly correlated (correlation coefficient of 0.88) which again shows that our agent managed to learn the desired reward function. The results in Fig. 6 are for the velocity experiment, the action norm scatter plot is in the appendix.

### v2

@@ 4.3 Curriculum Learning
To overcome this issue and avoid any serious degradation to task-specific performance, we devised a simple curriculum learning approach Bengio et al. (2009). Instead of updating the policy with the combined reward $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime})$, we use $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+\alpha(\bar{r}_{t}(s,a,s^{\prime}))f_{\theta}(s,a,s^{\prime})$ with the adaptive weighting $\alpha(\bar{r}_{t}(s,a,s^{\prime}))=\frac{\bar{r}_{ave}}{R_{MAX}}$ where $\bar{r}_{t}(s,a,s^{\prime})\in[0,R_{MAX}]$, and $\bar{r}_{ave}$ is the historical average of task-specific rewards over a window of 256 steps, for stability. This allows us to gradually increase the weighting of our learned reward as the policy improves on the main task. This way the common-sense reward does not have a significant effect at the beginning of training, and thus only impacts the policy after it had several update rounds. We note that this does not affect the discriminator update, which still uses Eq. 6. Also, when training a new agent with our learned reward, we use the standard aggregation $r_{t}(s,a,s^{\prime})=\bar{r}_{t}(s,a,s^{\prime})+f_{\theta}(s,a,s^{\prime})$.
4.4 Extension to Unknown Task Rewards
So far, we assumed that the task reward is known and focused on the task-agnostic reward. While we believe this is an important and common scenario, we will show our approach is not limited to this setting. Here, we extend our approach to the case where we have expert demonstrations from $T$ tasks, however the task-specific rewards are unknown. As before, we assume each expert maximizes a combination of the task-specific and common-sense behavior rewards, and we aim to recover a common-sense reward, $r_{cs}$, which will capture the shared desired behavior.
In these cases, we need to simultaneously learn per-task reward functions from each expert’s demonstrations and a shared common sense reward from all experts. For each expert, we train a task discriminator to learn the task-specific reward:
$D_{\phi_{t}}(s,a,s^{\prime})=\frac{\exp{(\bar{r}_{\phi_{t}}(s,a,s^{\prime}))}}{\exp{(\bar{r}_{\phi_{t}}(s,a,s^{\prime}))}+\pi_{w_{t}}(a|s)}.$ (8)
where $\phi_{t}$ are the learned parameters for the reward of task $t$. In addition, we train a shared discriminator for the common sense reward:
$D_{\theta}(s,a,s^{\prime})=\frac{\exp{(f_{\theta}(s,a,s^{\prime}))}}{\exp{(f_{\theta}(s,a,s^{\prime}))}+\pi_{w_{t}}(a|s)}.$ (9)
We then use each task reward and the shared common sense reward to update each task generator policy. This process of our extension for multi learned task (MT-CSIRL+LT) is depicted in Appendix A.4
@@ 5 Experiments
For our experiments, we use the Meta-world benchmark Yu et al. (2020b). It provides a diverse set of robotic manipulation tasks, which share the same robot, action space, and observation space. Thus, this benchmark is suitable for evaluating the effectiveness and transferability of our inverse reinforcement learning method across various scenarios. We trained all policies using the Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018). In order to emphasize our primary goal of learning transferable rewards, we limit our experimental setup to Meta-world tasks on which SAC performed well, according to the Meta-world’s paper. We used the exact training process and SAC hyperparameters as in Yu et al. (2020b). We repeat the experiments five times using different random initializations and report the mean and standard deviation of the performance. Using the Meta-world terminology, there are several distinct tasks, e.g., reach-wall and drawer-close. For each task, there are several variations with distinct targets. For example, in the drawer-close task, a different target would be a different location of the drawer. We also note that Meta-world has two versions for each task. All experiments in this paper use the second version, i.e., reach-wall is reach-wall-V2, and we omit the V2 for brevity.
@@
The goal of our experiments is to demonstrate that IRL fails to learn a useful cs-reward and that training the cs-reward on multiple tasks enables us to learn such a reward. To show that, we perform several experiments, each time learning the cs-reward (which we will describe shortly) from a more diverse set of tasks. See Fig. 8 in Appendix B for a visual summary of our experiments.
Common-Sense Rewards:
As Meta-world only contains task-specific rewards, we designed two simple common-sense rewards for our experiments. The explicit common-sense reward is not given directly to the agent during its IRL training process; instead, it was learned through expert demonstrations. The common-sense reward is used to score the agent, for evaluation purposes.
@@
Our first cs-reward directs the agent to move one of the key points along the robotic arm (whose 3D location is part of the observation vector) with a target velocity $v_{target}$. The reward function is given by
$r_{CS}(s,a,s^{\prime})=-C_{v}\cdot\lvert\|\ell(s^{\prime})-\ell(s))\|_{2}-v_{target}\rvert$, (10)
@@
where $\ell(s)$ is the 3D location of the selected key point given by the observation vector. The second reward directs the $L_{1}$ norm of the action towards a target value $n_{target}$. The reward is defined as follows,
$r_{CS}(s,a,s^{\prime})=-C_{n}\cdot|\|a\|_{1}-n_{target}|.$ (11)
@@
We name these reward functions the velocity and action norm cs-rewards, respectively. See Appendix B for further details. One important property of these rewards is that they do not depend on the task or environment and, as such, should be easily transferable. We specifically designed these simple, common-sense behaviors to show how IRL struggles to learn a transferable reward even in this simple setting.
Baselines:
In our experiments, we evaluate and compare the following methods: (1) Expert: RL trained with the task and ground-truth cs-reward; (2) SAC: An RL trained with only the task reward; (3) MT-AIRL Fu et al. (2018), Implementation of multi-task AIRL for learning the entire combined reward per task; (4) MT-VAIRL and (5) MT-VAIRL-GP Peng et al. (2018), implementation of multi-task VAIRL and multi-task VAIRL-GP for learning the entire combined reward per task. Our methods: (6) MT-CSIRL, (Alg. 1) and (7) MT-CSIRL+LT (Alg. 2). Since we introduce a novel setting, none of the existing standard solutions we compare to use the separation between task-agnostic reward and task-specific reward. However, they still allow us to infer the benefits of utilizing this assumption.
5.1 Learning CS-Reward from a Single Task
@@
We will show empirically that, even for simple common-sense rewards for which the IRL process successfully trains an agent, the final reward might not be useful for training a new agent from scratch. This issue occurs even when applying the reward on the exact same task and target. To show this, we train an expert with the task-specific (i.e., reach-wall) and ground truth common-sense reward for each of our common-sense rewards. In both velocity and action norm cases, we successfully train experts who maximize both the task rewards and the common-sense rewards. Then, for each expert, we sample a set of trajectories for IRL training. The cs-reward is learned using Alg. 1 with a single task (i.e., $T=1$). Finally, we train a new agent on the exact same task and target using our learned cs-reward and known task reward. In all of the following experiments in this section, all trained agents achieved 100% success rate on the main task. Therefore, we only report the results on the common-sense rewards.
@@
We present our results in Fig. 2 (a) & (b). During the IRL process, the trained agent effectively acquired the desired common-sense behavior, marked in the horizontal red line. However, when we train a new RL agent with our learned reward (and the task-specific reward) it is indistinguishable from the baseline SAC trained with only the task-specific reward. Similar results with different tasks are shown in Appendix A.1. We note that this phenomenon has been also previously observed in other bi-level optimization scenarios Navon et al. (2020); Vicol et al. (2022).
@@
In Fig. 2 (c), we also present a scatter plot of the learned cs-reward versus the ground-truth cs-reward. This shows a very small correlation between the learned reward and the ground-truth reward (correlation coefficient of 0.18), which further demonstrates that the IRL fails to capture the desired reward.
(a) Action norm experiment
(b) Velocity experiment
(c) Rewards’ correlation
Figure 2: Single task cs-reward: In (a), (b), we plot the rewards during the IRL process (IRL Agent), an RL agent with the learned cs-reward from the IRL process (Transferred), and a baseline RL agent without any common sense component (SAC). The red horizontal line represents the target value in velocity/action norm see Eq. 10 and 11.The scatter plot (c) shows the correlation between the ground-truth reward and the learned CS-Reward.
5.2 Learning CS-Reward from Multiple Targets
@@
In the next experiment we show that when we train our IRL on expert demonstration from different targets, we manage to learn a reward that can be used to train new agents on unseen targets. However, it does not transfer to novel tasks. Again, all agents trained in this section achieve 100% success rate so we only show results for the cs-reward. To do that, we train our MT-CSIRL method with experts trained on different targets for the same task, for each of our cs-rewards. We then test how these learned rewards can be used to train on new targets in this environment and how they transfer to other tasks. We execute this experiment on five different Meta-world tasks: reach-wall, button-press-topdown-wall (BP Topdown-wall), drawer-close, push-back and coffee-button. For each task, we train multiple experts, one for each target. We sample trajectories from the experts’ policies, and we use this trajectories as the expert demonstrations. We then learn a cs-reward using our method MT-CSIRL. We evaluate the performance of agents trained from scratch using this learned cs-reward on the training task as well as on new unseen tasks. For more experiment details see Appendix A.2.
@@
The results for action norm reward are presented in Table 1, and velocity results are in Appendix A.2. The numbers represent the ratio between the agents’ action norm and its target value, i.e., the closer to one, the better. On the diagonal, we see the results on different targets for the original task, with the baseline results for agents trained only on the task reward in parenthesis. On the off-diagonal, we see the results when transferring to new tasks. When looking at the diagonal elements of Tables 1, we see that our reward transfers well to new targets for the seen task. While our agents do not reach the target values, they still show significant improvement over the baseline. However, the drop in performance on the off-diagonal elements indicates that the learned reward does not transfer well to novel tasks. This is especially interesting when we consider our simple cs-rewards, which do not depend on the specifics of the environment (and independent of the state for action norm).
reach-wall BP Topdown-wall drawer-close push-back coffee-button
reach-wall $\mathbf{0.91}(0.57)$ $0.43$ $0.60$ $0.54$ $0.58$
BP Topdown-wall $0.60$ $\mathbf{0.91}(0.61)$ $0.58$ $0.60$ $0.59$
drawer-close $0.59$ $0.53$ $\mathbf{0.90}(0.56)$ $0.63$ $0.57$
push-back $0.56$ $0.56$ $0.51$ $\mathbf{0.90}(0.62)$ $0.55$
coffee-button $0.58$ $0.57$ $0.54$ $0.58$ $\mathbf{0.92}(0.48)$
Table 1: Action Norm cs-reward. Each row represents a different task for MT-CSIRL training, each column represents a task for RL training with learned reward and evaluation. The Numbers represent the ratio between the agents’ action norm and its target value. In the diagonal, Action Norm reward for training solely on the task reward appears in parentheses.
5.3 Learning CS-Reward from Multiple Tasks
@@
Here we show that when we train the IRL process on experts’ trajectories from multiple tasks, we learn a useful cs-reward that can be transferred to new unseen tasks. In order to do that, we perform our MT-CSIRL method again, but this time we train each expert on a different meta-world task.
@@
The results for velocity as cs-reward, and action norm as cs-reward are presented in table 2. As can be easily observed, the MT-CSIRL results show that our learned cs-reward manages to transfer the desired behavior even to unseen tasks, albeit not as strongly as the ground-truth reward. We included the MT-AIRL baseline, which unsurprisingly does not work well, to show the importance of our split between task-specific and task-independent rewards. Without this distinction combining different tasks is non-trivial and can easily harm performance as we see here. We also note that the MT-AIRL baseline shares similarities with Gleave and Habryka (2018), however, they train with a meta-learning approach similar to MAML.
@@
To further illustrate our results we show in Fig. 3 the ground truth cs-reward (both for velocity and action norm) on the unseen test task. The figures show the positive impact of our cs-reward, making the agent’s behavior significantly more aligned with the expert. Finally, we show in Fig. 3 a scatter plot of the ground-truth cs-reward versus our learned reward on a new unseen task. As one can see, these rewards are highly correlated (correlation coefficient of 0.88) which again shows that our agent managed to learn the desired reward function. The correlation results in Fig. 3 are for the velocity experiment, the action norm scatter plot is in Appendix A.3.
Action Norm Velocity
Agent Avg Action Norm Success-rate Avg velocity Success-rate
Expert $2.95\pm 0.01$ $100\pm 0.00$ $0.81\pm 0.02$ $100\pm 0.00$
SAC $1.57\pm 0.48$ $100\pm 0.00$ $0.32\pm 0.16$ $100\pm 0.00$
MT-AIRL $1.30\pm 0.47$ $63.5\pm 1.20$ $0.11\pm 0.01$ $67.5\pm 5.90$
MT-VAIRL $2.29\pm 0.07$ $79.6\pm 4.45$ $0.38\pm 0.06$ $73.3\pm 2.51$
MT-VAIRL-GP $2.32\pm 0.19$ $75.0\pm 1.37$ $0.39\pm 0.02$ $76.8\pm 2.65$
MT-CSIRL (Ours) $\mathbf{2.75}\pm\mathbf{0.18}$ $99.9\pm 0.01$ $\mathbf{0.73}\pm\mathbf{0.07}$ $99.2\pm 0.30$
MT-CSIRL+LT (Ours) $2.70\pm 0.35$ $97.9\pm 0.08$ $0.70\pm 0.04$ $97.2\pm 0.06$
Table 2: Experiment results on unseen tasks with action norm cs-reward and velocity cs-reward, both learned in a multi task learning framework. first with a given task reward during IRL (MT-CSIRL), and second with a learned task reward during IRL (MT-CSIRL+LT).
(a) Action norm CS reward
(b) Velocity CS reward
(c) Rewards’ correlation
Figure 3: Ground Truth CS-Reward: In (a) and (b), We show the ground-truth common sense behavior on three different experiments. Expert: maximizes task reward and ground-truth cs-reward. MT-CSIRL: trained with the learned cs-reward and with the task reward. SAC: "vanilla" training, trained only with task reward. In (c) we visualized the scatter plot between Ground-Truth & Learned CS-Reward from MT-CSIRL method, on Velocity Experiment.
5.4 Learning CS-Reward and Task-Reward from Multiple Tasks
In this section, we implement the extension to our methodology as detailed in 4.4, where we will assume we have expert demonstrations from T tasks but the task-specific rewards are unknown. We train an IRL process on multiple tasks using MT-CSIRL+LT method. Training this process gives as a learned cs-reward, and a learned task-specific reward for each task in the training process. In order to show that the learned cs-reward can be transferred to novel tasks, we train new RL agents with the ground truth task reward and our learned cs-reward (results in Table 2). The difference here is that during the cs-reward training we did not have access to the task rewards. In Appendix A.4 we show how this method performs on novel tasks without the ground truth reward but with expert demonstrations.